OpenStudy (anonymous):

Simplify the Expression, if possible x^2-x-12/x^2-6x+8 Simplify the Expression, if possible x^2+3x-10/x^2+5x+6

OpenStudy (anonymous):

(x+3)/(x-2)

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

oh maybe the first one sorry.

OpenStudy (anonymous):

second one top and bottom factor but nothing cancels

OpenStudy (anonymous):

which one are you talking about

OpenStudy (anonymous):

\[\frac{x^2+3x-10}{x^2+5x+6}\] \[=\frac{(x-2)(x+5)}{(x+2)(x+3)}\]

OpenStudy (anonymous):

jay, do you know how to factor quadratics?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

why dont you try factoring the quadratic on the numerator of the first problem?

OpenStudy (anonymous):

would it be 2 because all the numbers are even

OpenStudy (anonymous):

http://www.youtube.com/watch?v=eF6zYNzlZKQ

OpenStudy (anonymous):

ooooo but is it a easier way